Circular No. 8188

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003hp IN UGC 10942 Further to IAUC 8184, M. Moore and W. Li report the LOTOSS dis-

covery of an apparent supernova (mag ∼ 18.4) on unfiltered KAIT images taken on Aug. 25.2 and 27.2 UT. SN 2003hp is located at α = 17^{h}40^{m}18^{s}.69, δ = +51^{o}01′41′′.3 (equinox 2000.0), which is 17′′.2 east and 4′′.8 south of the nucleus of UGC 10942. A KAIT image taken on Aug. 17.2 UT showed nothing at this position (limiting mag ∼ 19.5).

ϵ INDI B

K. Volk, Gemini South Observatory (GSO), La Serena; R. Blum, Cerro Tololo Interamerican Observatory, La Serena; G. Walker, University of British Columbia; and P. Puxley, GSO, report their detection of a faint companion to the nearby brown dwarf ϵ Ind B (Scholz et al. 2003, A.Ap. 398, L29) on Aug. 18 and 20 with the Acquisition Camera and the Phoenix spectrometer at GSO. The separation is 0′′.62 ± 0′′.02 in p.a. 133^{o}.4 ± 0^{o}.5 in 1.28-μm narrowband images, corresponding to a projected separation of 2.1 AU at the distance of ϵ Ind B. Due to its large proper motion, ϵ Ind B is 17′′.6 from its discovery position, or now at α = 22^{h}04^{m}12^{s}.415, δ = −56^{o}47′05′′.66 (equinox 2000.0); there is nothing at this position in the 1999.855-epoch 2-MASS J-band image on which ϵ Ind B was discovered. The companion is bright enough to have been detected (at S/N = 300) in the 2-MASS image if it is a background star; this fact and its unusual colors lead the authors to tentatively identify it as a late T-type brown dwarf (cf. Burgasser et al. 2002, Ap.J. 564, 421; Geballe et al. 2002, Ap.J. 564, 466) or a large planet. The following relative brightnesses of the new object (as a percentage of the brightness of ϵ Ind B) were measured from I-band and various narrowband Phoenix filters: I ≈ 0.25; [1.083 μm], 0.53; [1.282 μm], 0.65; [1.556 μm], 0.29; [2.106 μm], 0.17; [2.321 μm], < 0.03.

PECULIAR VARIABLE IN CRUX W. Liller, Viña del Mar, Chile, reports the following magnitudes for

this star (cf. IAUC 8185): Aug. 22.008 UT, V = 9.91 ± 0.02 (CCD); 22.009, B = 11.16 ± 0.05 (CCD); 22.996, R = 9.3 ± 0.1 (red photographs).

Visual magnitude estimates: Aug. 22.905 UT, 10.3 (R. Y. Shida, São Paulo, Brazil); 23.931, 10.1 (J. G. S. Aguiar, Campinas, Brazil); 24.910, 10.0 (Aguiar); 25.542, 10.3 (A. Pearce, Nedlands, W. Australia); 26.510, 10.5 (Pearce).

2003 August 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green